OpenStudy (anonymous):

The Lopez family had dinner at a restaurant. The bill for the meal was $68 and they gave the waiter a 20% tip. What was the total cost of the dinner? A. $68.20 B. $70.00 C. $81.60 D. $88.00

OpenStudy (daniellelovee):

20% of 68

OpenStudy (ajspeller):

1.2 * 68

OpenStudy (daniellelovee):

68/100=x/20

OpenStudy (anonymous):

68 * .2 = 13.6 13.6 + 68 = ?

OpenStudy (ajspeller):

You pay 100% or the total cost of the dinner ... PLUS You pay 20% more for the tip 100% = 1 20% = 0.2 1+ 0.2 = 1.2 1.2 * 68 = Total cost

OpenStudy (domebotnos):

depends on how many people but i say 81 ish dollars

OpenStudy (anonymous):

c

OpenStudy (anonymous):

yes

OpenStudy (daniellelovee):

yup the answer is c

OpenStudy (domebotnos):

yay im right! :P lol

OpenStudy (anonymous):

who deserves the medal most?

OpenStudy (domebotnos):

ajspeller duh

OpenStudy (domebotnos):

lol